# MUST-RAG: MUSical Text Question Answering with Retrieval Augmented Generation

### v1

@@ 1 Introduction
Traditionally, domain adaptation of LLMs has often been achieved by fine-tuning them on domain-specific data [1, 2, 3]. However, this approach faces challenges in securing high-quality training data, and as model size increases, the training time and cost also rise significantly. Additionally, continuously updating the model with new knowledge remains a persistent challenge.
In this paper, we propose MusT-RAG, a framework that leverages Retrieval Augmented Generation (RAG) [4] techniques to enhance general-purpose LLMs for music-specific tasks. The core idea behind MusT-RAG is to augment LLMs with external knowledge retrieval mechanisms. Specifically, the model retrieves relevant external knowledge from a pre-constructed, comprehensive music-specific vector database in order to answer input questions.
For music-domain specific retrieval, we introduce MusWikiDB, which, to our knowledge, is the first comprehensively curated vector database designed specifically for music-related content. We explore various design choices for optimizing retrieval performance, including embedding models and chunking strategies. By incorporating this retrieval process, MusT-RAG enables LLMs to efficiently generate contextually relevant responses, drawing on specialized music knowledge to enhance performance on music-related tasks, all without requiring additional training. Furthermore, we extend the application of RAG beyond inference by incorporating contextual information during the fine-tuning process. Our empirical analysis reveals that this context-aware training enhances the model’s contextual understanding capabilities, defined as the ability to generate coherent and relevant text within a specific context [5], outperforming conventional fine-tuning approaches.
@@ 2 Music Question Answering
Question Answering (QA) refers to the task of providing an appropriate answer to a given question, which is one of the Information Retrieval (IR) tasks [6]. The task is typically framed as retrieving relevant information from a collection of documents or knowledge sources to answer fact-based questions. Open-domain QA involves answering questions from a vast and varied set of topics using a large collection of general knowledge documents [7, 8]. In contrast, domain-specific QA targets specialized fields such as medicine [9], law [10], or music [11, 12, 13, 14], where both the document set and the questions are confined to that domain. In this work, we define the MQA task as the problem of providing accurate and relevant answers to music-related questions by leveraging domain-specific musical knowledge.
Several recent studies have introduced music-related benchmarks to evaluate LLM performance. MuChoMusic [11] features 1,187 audio-based multiple-choice questions that assess both musical knowledge and reasoning capabilities. MusicTheoryBench [12] contains 372 expert-validated questions designed to evaluate advanced music knowledge and reasoning skills. TrustMus [13] comprises 400 questions across four domains—People, Instruments and Technology, Genres, Forms, and Theory, and Culture and History—all derived from The Grove Dictionary Online [15]. ZIQI-Eval [14] presents a comprehensive evaluation framework consisting of 14,000 comprehension tasks that span 10 major topics and 56 subtopics, encompassing a broad spectrum of music-related knowledge.
@@ 3.1.1 Indexing
Sparse Embeddings [16, 17] use term frequency-based scoring to match exact keywords, offering fast and interpretable retrieval for large-scale datasets.
Dense Embeddings [18, 19, 17] map questions and documents into a shared vector space, enabling semantic matching beyond keyword overlap.
Audio-Text Joint Embeddings [20, 21, 22, 23, 24] extend this further by jointly embedding text with the audio modality. By leveraging contrastive learning between audio and text, they can serve as more domain-specialized text embedding models for music-related tasks.
@@ 3.1.2 Retrieval
[…]
@@
[…]
@@ 3.1.3 Generation
LLMs often struggle with specialized tasks such as MQA due to limited exposure to domain-specific knowledge during pretraining. To address this, two primary domain adaptation strategies are commonly used: fine-tuning and RAG. Fine-tuning is akin to a closed-book exam: the model internalizes domain knowledge during training and must rely solely on that knowledge at inference. While effective for learning structured formats or stylistic patterns [26, 27], it is resource-intensive and inflexible when adapting to new or frequently changing knowledge. In contrast, RAG is like an open-book exam: the model dynamically retrieves relevant information from an external knowledge source during inference. This enables LLMs to access up-to-date and specialized information without retraining. Prior studies [28, 29] show that RAG improves factual accuracy, mitigates hallucinations, and provides greater transparency by allowing source verification. It is also more scalable and economically efficient, as it does not require updating model parameters [27]. These benefits are especially useful in rapidly evolving domains like music, where new artists, compositions, and styles continuously emerge.
@@ 4.1 MusWikiDB
Table 1 compares our proposed MusWikiDB with the Wikipedia corpus [8]. While MusWikiDB contains fewer pages (31K vs 3.2M) and has a smaller vocabulary size (786K vs 21.5M), it consists exclusively of music-specialized text information.
Based on the ablation study in Section  6.3, the text was then split into segments of up to 128 tokens, with a 10% overlap between adjacent passages, to preserve context between passages. For embedding, we employed BM25 [16], a classical and highly effective algorithm for ranking text relevance, which helped build an efficient index for MusWikiDB. This allowed us to quickly retrieve relevant information during RAG-based inference, improving the accuracy and relevance of answers. The resulting MusWikiDB provides a scalable, up-to-date knowledge base that enhances the performance of RAG in MQA tasks, allowing the system to answer complex, domain-specific music-related questions with more accuracy and context.
MusWikiDB Wikipedia Corpus [8]
@@
Table 1: MusWikiDB and Wikipedia Corpus [8] statistics.
@@ 4.2 ArtistMus
The existing text-only MQA benchmarks have focused on multimodal music understanding [11, 12] or musicology topics such as melody, chords, and history [12, 13]. However, there has been no benchmark that focuses on music metadata, particularly the artist, which is crucial in music listening contexts [30, 31]. Therefore, we created the ArtistMus to test the performance of LLMs in artist-related QA, using artist-related data from MusWikiDB.
We grouped sections into five categories: biography, career, discography, artistry, and collaborations. Token lengths ranging from 500 to 2000 were considered. Genre normalization [32] was applied by first converting all genre labels to lowercase, and then removing spaces, hyphens (-), and slashes (/). We obtained 48 root genres from  [33], and after retaining only the data corresponding to the top 300 most frequent genres, each genre was mapped to the 20 final genre labels. To extract artists’ regional information, we provided the abstract of pages to the Llama 3.1 8B Instruct [34] to extract information on the country of the artist. The country list was obtained from the pycountry library. Then, we select a diverse range of 500 artists based on topic, genre, and country. Country was set as the highest priority, with a preference for artists from minor countries. Subsequently, popular genres and topics were replaced with less common ones. We generated one factual and one contextual question for each artist to evaluate the LLM’s factuality and contextual understanding. To construct these questions, we provided GPT-4o [35] with the corresponding section text. Factual questions focus on verifiable details such as dates, names, or events, whereas contextual questions require reasoning or synthesis across multiple pieces of information within the passage.
@@ 5 Experiments
GPT-4o [35] N/A 70.0 64.8 67.4 93.2 92.8 93.0
ChatMusician [12] 7B 28.0 25.2 26.6 78.8 67.6 73.2
MuLLaMA [36] 7B 27.2 25.2 26.2 38.4 40.0 39.2
Llama 3.1 8B Instruct [34] 8B 40.0 38.0 39.0 87.6 82.8 85.2
@@
GPT-4o [35] N/A 48.0 47.0 57.0 60.0 53.0
ChatMusician [12] 7B 18.0 20.0 26.0 24.0 20.0
MuLLaMA [36] 7B 25.0 15.0 18.0 21.0 19.8
Llama 3.1 8B Instruct [34] 8B 36.0 24.0 41.0 42.0 35.8
@@ 5.2 Models
We compare zero-shot and QA fine-tuned models with our proposed RAG inference and RAG fine-tuned models to evaluate MQA performance. Following [11], we consider a response incorrect if it deviates from the expected format.
Zero-shot Baselines   We evaluated GPT-4o [35] (API-based), Llama 3.1 8B Instruct [34] (open-source), and two music-specific models: MuLLaMA [36] and ChatMusician [12]. MuLLaMA is designed to handle audio based question answering. ChatMusician specializes in music understanding and generation with ABC notation.
QA Fine-tuning   We fine-tune the Llama 3.1 8B Instruct [34] on 8K multiple-choice QA pairs that were generated from MusWikiDB.
RAG Inference   We use Llama 3.1 8B Instruct [34] as our base model and implement RAG at inference-time using MusWikiDB as the retrieval database.
@@ 5.3 Training Configurations
The models are trained for one epoch using LoRA [37] with 8-bit quantization with the following hyperparameter settings: batch size = 2, gradient accumulation steps = 4, learning rate = 3e-5, weight decay = 0.005, warmup ratio = 0.1, cosine scheduler [38], AdamW [39] optimizer, r = 16, alpha = 16, and dropout = 0.1. For the ArtistMus dataset, half of the artists were included in the training data (Seen), while the other half were excluded (Unseen).
@@ 5.4 Retriever Configurations
To select the optimal retriever configuration MusWikiDB, we performed an ablation study using the ArtistMus benchmark. We varied the passage size (128, 256, 512 tokens) and embedding models (BM25 [16], Contriever [19], CLAP [20]). For CLAP, we increased the token limit without additional training. To ensure a fair comparison, we constrained the total token budget to 1024 by adjusting the number of retrieved passages: top-8 for 128-token passages, top-4 for 256, and top-2 for 512.
@@ 6.1 In-domain Performance
Zero-shot Baselines   As shown in Table 2, all models performed significantly worse on factual questions than on contextual ones, indicating challenges in recalling concrete information such as names or dates. GPT-4o [35] outperformed Llama [34] by 28.4% in factual performance, though the gap narrowed to 7.8% for contextual understanding. Despite being music-specific, both ChatMusician [12] and MuLLaMA [36] showed relatively low performance. ChatMusician slightly underperformed compared to Llama, while MuLLaMA exhibited the lowest scores, likely due to its lack of training on the MQA task and poor instruction-following capabilities.
@@
RAG Inference   By utilizing RAG inference without additional training, we were able to address the low factual performance that was an issue with previous LLMs. It demonstrated a 14.6% higher factual performance compared to GPT-4o [35]. Contextual performance improved by 3.6% compared to zero-shot, but was still 4.2% lower than GPT-4o.
@@ 6.2 Out-of-domain Performance
To validate the effectiveness of the MusT-RAG in out-of-domain scenarios, we conducted experiments by changing the benchmark from ArtistMus to TrustMus [13], using the same framework with in-domain evaluation. The results are presented in Table  3.
Zero-shot Baselines   A similar trend was observed in the zero-shot evaluation for the in-domain setting. MuLLaMA [36] and ChatMusician [12] performed worse than the random baseline (25%), which is due to incorrect answers being counted when the models failed to follow instructions. Given that the overall zero-shot performance closely aligns with the factual scores from the in-domain evaluation, we infer that TrustMus mostly consists of factual questions. The Llama 3.1 8B Instruct [34] model scored 17.2% lower than GPT-4o [35].
@@
RAG Inference   RAG inference led to an 5.0% performance improvement over zero-shot, demonstrating that MusT-RAG framework is also helpful for out-of-domain data, such as The Grove Dictionary Online [15], which is the basis for the TrustMus benchmark.
@@
Table 4: Llama 3.1 8B Instruct [34] RAG performance on ArtistMus, by different passage size and embeddings.
@@ 6.3 Ablation Study: Retriever Configurations
Table 4 shows the results of the RAG inference for the Llama 3.1 8B Instruct [34] with various passage sizes and embeddings, evaluated under the same total computation budget for fair comparison. The performance on contextual questions tended to improve as the passage size decreased across all embedding models. In contrast, for factual questions, only Contriever [19] showed clear improvements with shorter passages, while BM25 [16] and CLAP [20] showed little to no change in performance across different passage lengths. For factual questions, there was a significant performance gap between BM25 and the other two dense embeddings. This is likely because ArtistMus places high importance on music entities such as artist and albums. Overall, the best performance was achieved using BM25 with a passage size of 128. When compared to the gold context, the factual performance was 15.6% lower, and the contextual performance was 8.0% lower. In Figure 2, we compare the RAG inference performance using the Wikipedia corpus [8] and MusWikiDB. The results show that MusWikiDB achieves a 10x faster retrieval speed and 5.9% higher performance.
Figure 2: RAG performance and retrieval time for Wikipedia Corpus [8] and MusWikiDB.

### v2

@@ 1 Introduction
Traditionally, domain adaptation of LLMs has often been achieved by fine-tuning them on domain-specific data [jeong2024fine, sahoo2024large, satterfield2024fine]. However, this approach faces challenges in securing high-quality training data, and as model size increases, the training time and cost also rise significantly. Additionally, continuously updating the model with new knowledge remains a persistent challenge.
In this paper, we propose MusT-RAG, a framework that leverages Retrieval Augmented Generation (RAG) [lewis2021rag] techniques to enhance general-purpose LLMs for music-specific tasks. The core idea behind MusT-RAG is to augment LLMs with external knowledge retrieval mechanisms. Specifically, the model retrieves relevant external knowledge from a pre-constructed, comprehensive music-specific vector database in order to answer input questions.
For music-domain specific retrieval, we introduce MusWikiDB, which, to our knowledge, is the first comprehensively curated vector database designed specifically for music-related content. We explore various design choices for optimizing retrieval performance, including embedding models and chunking strategies. By incorporating this retrieval process, MusT-RAG enables LLMs to efficiently generate contextually relevant responses, drawing on specialized music knowledge to enhance performance on music-related tasks, all without requiring additional training. Furthermore, we extend the application of RAG beyond inference by incorporating contextual information during the fine-tuning process. Our empirical analysis reveals that this context-aware training enhances the model’s contextual understanding capabilities, defined as the ability to generate coherent and relevant text within a specific context [zhao2024enhancing], outperforming conventional fine-tuning approaches.
@@ 2 Music Question Answering
Question Answering (QA) refers to the task of providing an appropriate answer to a given question, which is one of the Information Retrieval (IR) tasks [allam2012question]. The task is typically framed as retrieving relevant information from a collection of documents or knowledge sources to answer fact-based questions. Open-domain QA involves answering questions from a vast and varied set of topics using a large collection of general knowledge documents [rajpurkar2016squad, karpukhin2020dense]. In contrast, domain-specific QA targets specialized fields such as medicine [pal2022medmcqa], law [chalkidis2021lexglue], or music [weck2024muchomusic, yuan2024chatmusician, ramoneda2024role, li2024music], where both the document set and the questions are confined to that domain. In this work, we define the MQA task as the problem of providing accurate and relevant answers to music-related questions by leveraging domain-specific musical knowledge.
Several recent studies have introduced music-related benchmarks to evaluate LLM performance. MuChoMusic [weck2024muchomusic] features 1,187 audio-based multiple-choice questions that assess both musical knowledge and reasoning capabilities. MusicTheoryBench [yuan2024chatmusician] contains 372 expert-validated questions designed to evaluate advanced music knowledge and reasoning skills. TrustMus [ramoneda2024role] comprises 400 questions across four domains—People, Instruments and Technology, Genres, Forms, and Theory, and Culture and History—all derived from The Grove Dictionary Online [sadie2001new]. ZIQI-Eval [li2024music] presents a comprehensive evaluation framework consisting of 14,000 comprehension tasks that span 10 major topics and 56 subtopics, encompassing a broad spectrum of music-related knowledge.
@@ 3.1.1 Indexing
Sparse Embeddings [robertson1994some, behnamghader2024llm2vec] use term frequency-based scoring to match exact keywords, offering fast and interpretable retrieval for large-scale datasets.
Dense Embeddings [devlin2019bert, izacard2021unsupervised, behnamghader2024llm2vec] map questions and documents into a shared vector space, enabling semantic matching beyond keyword overlap.
Audio-Text Joint Embeddings [wu2023large, manco2022contrastive, doh2023toward, doh2024enriching, wu2025clamp] extend this further by jointly embedding text with the audio modality. By leveraging contrastive learning between audio and text, they can serve as more domain-specialized text embedding models for music-related tasks.
@@ 3.1.2 Retrieval
where $q$ is the input question, $D$ is the entire database of text passages, and $c\subset D$ is the filtered context consisting of the top-$k$ passages, such that $|c|=k\ll|D|$. Each passage  [^1] A passage refers to a portion of a document relevant to a query [wade2005passage]. $p\in D$ is scored based on its similarity to the input question using cosine similarity between their embeddings:
@@
Here, $E(\cdot)$ denotes an embedding function that maps both questions and passages into a shared vector space. The retriever ranks all passages in $D$ by their similarity scores and selects the top-$k$ passages to form $c$, which serve as the external context for the generation step.
@@ 3.1.3 Generation
LLMs often struggle with specialized tasks such as MQA due to limited exposure to domain-specific knowledge during pretraining. To address this, two primary domain adaptation strategies are commonly used: fine-tuning and RAG. Fine-tuning is akin to a closed-book exam: the model internalizes domain knowledge during training and must rely solely on that knowledge at inference. While effective for learning structured formats or stylistic patterns [shuster2021retrieval, borgeaud2022improving], it is resource-intensive and inflexible when adapting to new or frequently changing knowledge. In contrast, RAG is like an open-book exam: the model dynamically retrieves relevant information from an external knowledge source during inference. This enables LLMs to access up-to-date and specialized information without retraining. Prior studies [yasunaga2022deep, wang2023self] show that RAG improves factual accuracy, mitigates hallucinations, and provides greater transparency by allowing source verification. It is also more scalable and economically efficient, as it does not require updating model parameters [borgeaud2022improving]. These benefits are especially useful in rapidly evolving domains like music, where new artists, compositions, and styles continuously emerge.
@@ 4.1 MusWikiDB
Table 1 compares our proposed MusWikiDB with the Wikipedia corpus [karpukhin2020dense]. While MusWikiDB contains fewer pages (31K vs 3.2M) and has a smaller vocabulary size (786K vs 21.5M), it consists exclusively of music-specialized text information.
Based on the ablation study in Section  6.3, the text was then split into segments of up to 128 tokens, with a 10% overlap between adjacent passages, to preserve context between passages. For embedding, we employed BM25 [robertson1994some], a classical and highly effective algorithm for ranking text relevance, which helped build an efficient index for MusWikiDB. This allowed us to quickly retrieve relevant information during RAG-based inference, improving the accuracy and relevance of answers. The resulting MusWikiDB provides a scalable, up-to-date knowledge base that enhances the performance of RAG in MQA tasks, allowing the system to answer complex, domain-specific music-related questions with more accuracy and context.
MusWikiDB Wikipedia Corpus [karpukhin2020dense]
@@
Table 1: MusWikiDB and Wikipedia Corpus [karpukhin2020dense] statistics.
@@ 4.2 ArtistMus
The existing text-only MQA benchmarks have focused on multimodal music understanding [weck2024muchomusic, yuan2024chatmusician] or musicology topics such as melody, chords, and history [yuan2024chatmusician, ramoneda2024role]. However, there has been no benchmark that focuses on music metadata, particularly the artist, which is crucial in music listening contexts [lee2010analysis, doh2024music]. Therefore, we created the ArtistMus to test the performance of LLMs in artist-related QA, using artist-related data from MusWikiDB.
We grouped sections into five categories: biography, career, discography, artistry, and collaborations. Token lengths ranging from 500 to 2000 were considered. Genre normalization [schreiber2015improving] was applied by first converting all genre labels to lowercase, and then removing spaces, hyphens (-), and slashes (/). We obtained 48 root genres from  [schreiber2016genre], and after retaining only the data corresponding to the top 300 most frequent genres, each genre was mapped to the 20 final genre labels. To extract artists’ regional information, we provided the abstract of pages to the Llama 3.1 8B Instruct [grattafiori2024llama] to extract information on the country of the artist. The country list was obtained from the pycountry library. Then, we select a diverse range of 500 artists based on topic, genre, and country. Country was set as the highest priority, with a preference for artists from minor countries. Subsequently, popular genres and topics were replaced with less common ones. We generated one factual and one contextual question for each artist to evaluate the LLM’s factuality and contextual understanding. To construct these questions, we provided GPT-4o [achiam2023gpt] with the corresponding section text. Factual questions focus on verifiable details such as dates, names, or events, whereas contextual questions require reasoning or synthesis across multiple pieces of information within the passage.
@@ 5 Experiments
GPT-4o [achiam2023gpt] N/A 70.0 64.8 67.4 93.2 92.8 93.0
ChatMusician [yuan2024chatmusician] 7B 28.0 25.2 26.6 78.8 67.6 73.2
MuLLaMA [liu2024music] 7B 27.2 25.2 26.2 38.4 40.0 39.2
Llama 3.1 8B Instruct [grattafiori2024llama] 8B 40.0 38.0 39.0 87.6 82.8 85.2
@@
GPT-4o [achiam2023gpt] N/A 48.0 47.0 57.0 60.0 53.0
ChatMusician [yuan2024chatmusician] 7B 18.0 20.0 26.0 24.0 20.0
MuLLaMA [liu2024music] 7B 25.0 15.0 18.0 21.0 19.8
Llama 3.1 8B Instruct [grattafiori2024llama] 8B 36.0 24.0 41.0 42.0 35.8
@@ 5.2 Models
We compare zero-shot and QA fine-tuned models with our proposed RAG inference and RAG fine-tuned models to evaluate MQA performance. Following [weck2024muchomusic], we consider a response incorrect if it deviates from the expected format.
Zero-shot Baselines   We evaluated GPT-4o [achiam2023gpt] (API-based), Llama 3.1 8B Instruct [grattafiori2024llama] (open-source), and two music-specific models: MuLLaMA [liu2024music] and ChatMusician [yuan2024chatmusician]. MuLLaMA is designed to handle audio based question answering. ChatMusician specializes in music understanding and generation with ABC notation.
QA Fine-tuning   We fine-tune the Llama 3.1 8B Instruct [grattafiori2024llama] on 8K multiple-choice QA pairs that were generated from MusWikiDB.
RAG Inference   We use Llama 3.1 8B Instruct [grattafiori2024llama] as our base model and implement RAG at inference-time using MusWikiDB as the retrieval database.
@@ 5.3 Training Configurations
The models are trained for one epoch using LoRA [hu2022lora] with 8-bit quantization with the following hyperparameter settings: batch size = 2, gradient accumulation steps = 4, learning rate = 3e-5, weight decay = 0.005, warmup ratio = 0.1, cosine scheduler [wolf-etal-2020-transformers], AdamW [loshchilov2017decoupled] optimizer, r = 16, alpha = 16, and dropout = 0.1. For the ArtistMus dataset, half of the artists were included in the training data (Seen), while the other half were excluded (Unseen).
@@ 5.4 Retriever Configurations
To select the optimal retriever configuration MusWikiDB, we performed an ablation study using the ArtistMus benchmark. We varied the passage size (128, 256, 512 tokens) and embedding models (BM25 [robertson1994some], Contriever [izacard2021unsupervised], CLAP [wu2023large]). For CLAP, we increased the token limit without additional training. To ensure a fair comparison, we constrained the total token budget to 1024 by adjusting the number of retrieved passages: top-8 for 128-token passages, top-4 for 256, and top-2 for 512.
@@ 6.1 In-domain Performance
Zero-shot Baselines   As shown in Table 2, all models performed significantly worse on factual questions than on contextual ones, indicating challenges in recalling concrete information such as names or dates. GPT-4o [achiam2023gpt] outperformed Llama [grattafiori2024llama] by 28.4% in factual performance, though the gap narrowed to 7.8% for contextual understanding. Despite being music-specific, both ChatMusician [yuan2024chatmusician] and MuLLaMA [liu2024music] showed relatively low performance. ChatMusician slightly underperformed compared to Llama, while MuLLaMA exhibited the lowest scores, likely due to its lack of training on the MQA task and poor instruction-following capabilities.
@@
RAG Inference   By utilizing RAG inference without additional training, we were able to address the low factual performance that was an issue with previous LLMs. It demonstrated a 14.6% higher factual performance compared to GPT-4o [achiam2023gpt]. Contextual performance improved by 3.6% compared to zero-shot, but was still 4.2% lower than GPT-4o.
@@ 6.2 Out-of-domain Performance
To validate the effectiveness of the MusT-RAG in out-of-domain scenarios, we conducted experiments by changing the benchmark from ArtistMus to TrustMus [ramoneda2024role], using the same framework with in-domain evaluation. The results are presented in Table  3.
Zero-shot Baselines   A similar trend was observed in the zero-shot evaluation for the in-domain setting. MuLLaMA [liu2024music] and ChatMusician [yuan2024chatmusician] performed worse than the random baseline (25%), which is due to incorrect answers being counted when the models failed to follow instructions. Given that the overall zero-shot performance closely aligns with the factual scores from the in-domain evaluation, we infer that TrustMus mostly consists of factual questions. The Llama 3.1 8B Instruct [grattafiori2024llama] model scored 17.2% lower than GPT-4o [achiam2023gpt].
@@
RAG Inference   RAG inference led to an 5.0% performance improvement over zero-shot, demonstrating that MusT-RAG framework is also helpful for out-of-domain data, such as The Grove Dictionary Online [sadie2001new], which is the basis for the TrustMus benchmark.
@@
Table 4: Llama 3.1 8B Instruct [grattafiori2024llama] RAG performance on ArtistMus, by different passage size and embeddings.
@@ 6.3 Ablation Study: Retriever Configurations
Table 4 shows the results of the RAG inference for the Llama 3.1 8B Instruct [grattafiori2024llama] with various passage sizes and embeddings, evaluated under the same total computation budget for fair comparison. The performance on contextual questions tended to improve as the passage size decreased across all embedding models. In contrast, for factual questions, only Contriever [izacard2021unsupervised] showed clear improvements with shorter passages, while BM25 [robertson1994some] and CLAP [wu2023large] showed little to no change in performance across different passage lengths. For factual questions, there was a significant performance gap between BM25 and the other two dense embeddings. This is likely because ArtistMus places high importance on music entities such as artist and albums. Overall, the best performance was achieved using BM25 with a passage size of 128. When compared to the gold context, the factual performance was 15.6% lower, and the contextual performance was 8.0% lower. In Figure 2, we compare the RAG inference performance using the Wikipedia corpus [karpukhin2020dense] and MusWikiDB. The results show that MusWikiDB achieves a 10x faster retrieval speed and 5.9% higher performance.
Figure 2: RAG performance and retrieval time for Wikipedia Corpus [karpukhin2020dense] and MusWikiDB.